You will receive $100 from a zero-coupon savings bond in 3 years. The nominal interest rate is 7.80%.

a. What is the present value of the proceeds from the bond? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Present Value:
b. If the inflation rate over the next few years is expected to be 2.80%, what will the real value of the $100 payoff be in terms of today’s dollars? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Real Value:
c. What is the real interest rate? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Real Interest Rate (In percent):
d. Show that the real payoff from the bond [from part (b)] discounted at the real interest rate [from part (c)] gives the same present value for the bond as you found in part (a). (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Present Value:
Answer #1

1.
=Future Value/(1+nominal rate)^t
=100/1.078^3
=79.825880

2.
=Future Value/(1+inflation rate)^t
=100/1.028^3
=92.04933514

3.
=(1+nominal rate)/(1+inflation rate)-1
=1.078/1.028-1
=4.86%

4.
=92.04933514/(1+1.078/1.028-1)^3
=79.825880